# Leveraging Offline Data in Linear Latent Bandits

### v1

@@ E.1 Confidence bound for the low-dimensional reward parameter
The rest of the proof is similar to Theorem 2 in Abbasi-Yadkori et al. [2011]. We first note that for any $x$, we have that
@@
Now note that $\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}\leq\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{2}\sqrt{\mu}\leq R\sqrt{\mu}$ and by the self normalized martingale concentration inequality from Abbasi-Yadkori et al. [2011] applied to $d_{K}$ dimensional vectors $\hat{\textbf{U}}^{\top}\textbf{X}_{t}$, we have that with probability at least $1-\delta$, $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\leq CR\sqrt{d_{A}\log(t/\delta)}$ for some universal constant $C$. So we have with probability at least $1-\delta$ that
@@
The rest of the proof is similar to Theorem 2 in Abbasi-Yadkori et al. [2011]. We first note that for any $x$, we have that
@@
[…]
Now note that $\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}\leq\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{2}\sqrt{\mu}\leq R\sqrt{\mu}$ and by the self normalized martingale concentration inequality from Abbasi-Yadkori et al. [2011] applied to $d_{K}$ dimensional vectors $\hat{\textbf{U}}^{\top}\textbf{X}_{t}$, we have that with probability at least $1-\delta$, $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\leq CR\sqrt{d_{A}\log(t/\delta)}$ for some universal constant $C$. Also recall that $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{\textbf{V}_{1,t}^{-1}}=\kappa_{t}$. So we have with probability at least $1-\delta$ that
@@
This follows immediately from Lemma 11 of Abbasi-Yadkori et al. [2011]. ∎
@@ E.2 Proof of the theorem
If our features are isotropic, then this is only maximized by a feature in the span of $\hat{\textbf{U}}$. So, the conditions of the first bound in Lemma 6 are fulfilled. We will denote $\textbf{V}_{1,t}=\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}}$. We have that with probability $21-\delta/$ 2, for all $x,a,t$, the following holds.
@@
Bounding regret when $\tau\Delta_{\operatorname{off}}\left(\sqrt{T}+\sqrt{\frac{d_{K}}{T}\sum_{t=1}^{T}\kappa_{t}^{2}}\right)\geq d_{A}$: In this regime, after the first timestep $S$ satisfying $\Delta_{\operatorname{off}}\left(\sqrt{S}+\sqrt{\frac{d_{K}}{S}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\geq d_{A}$, we run standard LinUCB with dimension $d_{A}$ and incur $\widetilde{O}(d_{A}\sqrt{T})$ regret with probability $1-\delta$. Until timestep $S-1$, we run the projected and modified version of LinUCB and incur regret bounded by

### v2

@@ E.1 Confidence bound for the low-dimensional reward parameter
The rest of the proof is similar to Theorem 2 in Abbasi-Yadkori et al., (2011). We first note that for any $x$, we have that
@@
Now note that $\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}\leq\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{2}\sqrt{\mu}\leq R\sqrt{\mu}$ and by the self normalized martingale concentration inequality from Abbasi-Yadkori et al., (2011) applied to $d_{K}$ dimensional vectors $\hat{\textbf{U}}^{\top}\textbf{X}_{t}$, we have that with probability at least $1-\delta$, $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\leq CR\sqrt{d_{A}\log(t/\delta)}$ for some universal constant $C$. So we have with probability at least $1-\delta$ that
@@
The rest of the proof is similar to Theorem 2 in Abbasi-Yadkori et al., (2011). We first note that for any $x$, we have that
@@
Here, $(i)$ holds because for any matrices $\textbf{A},\textbf{C}$ and any vector v, we have that $\|\textbf{A}\textbf{v}\|_{\textbf{C}}=\sqrt{\textbf{v}^{\top}\textbf{A}^{\top}\textbf{C}\textbf{A}\textbf{v}}\leq\|\textbf{v}\|_{2}\sqrt{\|\textbf{A}^{\top}\textbf{C}\textbf{A}\|_{2}}=\|\textbf{v}\|_{2}\|\textbf{A}\|_{\textbf{C}}$, where we recall that $\|\textbf{A}\|_{\textbf{C}}:=\sqrt{\|\textbf{A}^{\top}\textbf{C}\textbf{A}\|_{2}}$.
Now note that $\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}\leq\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{2}\sqrt{\mu}\leq R\sqrt{\mu}$ and by the self normalized martingale concentration inequality from Abbasi-Yadkori et al., (2011) applied to $d_{K}$ dimensional vectors $\hat{\textbf{U}}^{\top}\textbf{X}_{t}$, we have that with probability at least $1-\delta$, $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\leq CR\sqrt{d_{A}\log(t/\delta)}$ for some universal constant $C$. Also recall that $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{\textbf{V}_{1,t}^{-1}}=\kappa_{t}$. So we have with probability at least $1-\delta$ that
@@
This follows immediately from Lemma 11 of Abbasi-Yadkori et al., (2011). ∎
@@ E.2 Proof of the theorem
If our features are isotropic, then this is only maximized by a feature in the span of $\hat{\textbf{U}}$. So, the conditions of the first bound in Lemma 6 are fulfilled. We will denote $\textbf{V}_{1,t}=\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}}$. We have that with probability $1-\delta/2$, for all $x,a,t$, the following holds.
@@
Bounding regret when $\tau\Delta_{\operatorname{off}}\left(\sqrt{T}+\sqrt{\frac{d_{K}}{T}\sum_{t=1}^{T}\kappa_{t}^{2}}\right)\geq d_{A}$: In this regime, after the first timestep $S$ satisfying $\Delta_{\operatorname{off}}\left(\sqrt{S}+\sqrt{\frac{d_{K}}{S}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\geq d_{A}$, we run standard LinUCB with dimension $d_{A}$ and incur $\widetilde{O}(d_{A}\sqrt{T})$ regret with probability $1-\delta$. Until timestep S1S-1, we run the projected and modified version of LinUCB and incur regret bounded by